a) Given a utility function





u = x_{1} ^ 0.5 * x_{2} ^ 0.25





If P_{1} = 2 and P_{2} = 4 and Y=100, compute optimal X_{1} and X_{2} that will maximize utility

1
Expert's answer
2022-02-14T13:41:26-0500

"U= X_1^{0.5}X_2^{0.25}"

For utility maximization,

"\frac{Mu_1}{Mu_2}= \frac{P_1}{P_2}"

"Mu_1= 0.5X_1^{-0.5}X_2^{0.25}"

"Mu_2= 0.25X_1^{0.5}X_2^{-0.75}"

"\frac{0.5 X_ 2}{0.25X_1}= \frac{2}{4}"


"0.5X_1= 2X_2"

"X_1= 4X_2"


"X_2= 0.5X_1"

Budget line

"2X_1+4X_2= 100"


"2(4X_2)+ 4X_2= 100"

"X_2= 8.33"


"2X_1+ 4(0.5X_1)= 100"

"X_1= \frac{100}{4}= 25"